### v3

@@ Abstract
We consider estimation and inference of the effects of a policy in the absence of a control group. We obtain unbiased estimators of individual (heterogeneous) treatment effects and a consistent and asymptotically normal estimator of the average treatment effect. Our estimator averages over unbiased forecasts of individual counterfactuals, based on a (short) time series of pre-treatment data. The paper emphasizes the importance of focusing on forecast unbiasedness rather than accuracy when the end goal is estimation of average treatment effects. We show that simple basis function regressions ensure forecast unbiasedness for a broad class of data-generating processes for the counterfactuals, even in short panels. In contrast, model-based forecasting requires stronger assumptions and is prone to misspecification and estimation bias. We show that our method can replicate the findings of some previous empirical studies, but without using a control group.
@@ 1 Introduction
Evaluating policies, or treatments, is a central task in causal inference. Such evaluations usually require the construction of counterfactuals, i.e., outcomes that the treated would have experienced had they not received the treatment. When evaluating policies with partial participation, the convention is to construct counterfactuals for the treated using observed data on the control, i.e. on those who do not receive the treatment and who are observed at every time period as the treated. When there is universal participation, [^3] Some examples of universal policies: Medicare (e.g., Sun and Shapiro (2022)), child care programs (e.g., Baker et al. (2008)), mandatory job search programs (e.g., Blundell et al. (2004)), environmental policies (e.g., Gallego et al. (2013)), country-level trade agreements (Trefler (2004)), fiscal policies (Jorda and Taylor (2016)), state-level texting bans (Abouk and Adams (2013)), global pandemics such as COVID-19. researchers in economics typically rely on structural models for policy evaluation. [^4] For example, Heckman and Vytlacil (2005) discuss how functional form restrictions and support conditions can be used to substitute for the lack of control individuals.
In this paper, we introduce an alternative approach for the evaluation of policies with universal participation. While our approach does not rely on structural models, it is unique in its focus on understanding what assumptions on the data-generating process (DGP) for the counterfactuals are needed to consistently estimate the effect of a universal treatment. With partial participation, the literature has not conventionally discussed the DGP, opting instead for high-level assumptions such as unconfoundedness or parallel paths. Recent literature has illustrated the types of assumptions implicitly imposed on individual dynamic choice by these high-level assumptions (e.g., Ghanem et al. (2022); Marx et al. (2023)), and our approach can be seen as reflecting a possibly growing interest in making assumptions on the DGP explicit.  [^5] Additionally, our approach can also be used as a robustness check when there is a control group, but its validity is uncertain or the assumptions that existing methods impose may be too strong to justify. For example, estimators for the treatment effect under unconfoundedness require that researchers choose both which observed covariates to condition on to guarantee that unconfoundedness holds (and include them in the propensity score) and the functional form of the propensity score. Both are fraught with specification uncertainty, e.g., Hirano and Imbens (2001); Kitagawa and Muris (2016). Likewise, difference-in-differences-type estimators require some sort of parallel-paths assumption. Although there is work that allows for a variety of robustness and sensitivity analyses, researchers must choose how to construct both the control group and the weights placed on different time periods, see, e.g., Roth et al. (2023).
Our parameter of interest is the average treatment effect on the treated (ATT). [^6] With universal participation, the ATT is the same as the average treatment effect – the ATE. Our baseline method uses individual time series of pre-treatment outcomes to forecast individual counterfactuals. The time series can be short. The cross-sectional average of individual treatment effects, i.e. the individual differences between the observed post-treatment outcome at a particular time and the forecasted counterfactual, is taken as an estimate of the ATT at that time period. We call our estimator Forecasted Average Treatment effect (FAT). We show that, even in short panels, FAT is a consistent and asymptotically normal estimator of the ATT under two high-level assumptions: (1) the forecasts are unbiased on average, and (2) the average differences between the observed post-treated outcomes and the forecasted counterfactuals satisfy a central limit theorem. The main contribution of this paper is to characterize both the class of DGPs for the individual counterfactuals and the forecast methods that satisfy these requirements.
Focusing on individual-level forecasts allows for heterogeneous treatment effects, unbalanced panels, and staggered (or individual-specific) treatment timing. On the other hand, achieving robustness to the absence of a valid control group naturally comes at a cost: Our baseline method rules out unforecastable common shocks that affect all treated individuals between the time of implementation and evaluation of the treatment. In the presence of unforecastable common shocks, recovering the ATT requires access to a (control) group that is subject to the shock but not to the treatment, see Section C.2 in the Online Appendix. [^7] Alternatively, it requires access to another outcome variable that is subject to the common shock but not to the treatment, see Section C.1 in the Online Appendix. This illustrates the trade-offs involved in recovering the ATT: Assumptions must be placed on either a control group if available, or on the counterfactuals of the treated group. A key insight of this paper is that, surprisingly, the assumptions on the counterfactuals of the treated group do not include correct model specification. All that is needed are forecasts of the counterfactuals that are unbiased on average across individuals.
Forecast unbiasedness is typically not the main concern in time series and panel forecasting, which focuses instead on accuracy (i.e., variance reduction at the cost of some bias). Our first contribution is to show that forecasts based on basis function regressions, i.e. forecasts based on regressing individual pre-treatment outcomes on known basis functions of time (such as polynomial time trends), deliver unbiased estimators of the individual treatment effects, and that FAT based on these forecasts is a consistent and asymptotically normal estimator of the ATT. Our second contribution is to show that forecast unbiasedness holds for a large class of DGPs that expresses the counterfactuals as, potentially, the sum of two individual-specific unobserved components: a mean-stationary process and a random walk. Importantly, this is a large and flexible class that encompasses, e.g., linear panel models with fixed effects, lagged outcomes with heterogeneous autoregressive parameters, and stochastic trends. [^8] This is a much larger class of models than can currently be accommodated by methods in the panel literature with short-T, where practitioners must restrict their analyses to stationary processes, and must choose between fixed effects and lagged outcomes (with homogeneous autoregressive parameters), e.g., Angrist and Pischke (2009). If the DGP includes a deterministic trend, an additional requirement is imposed. The additional requirement is that the basis functions used to produce the forecast must be correctly specified. This result emphasizes that it is not necessary to model the stochastic component of the counterfactuals to obtain unbiased forecasts. Rather, the focus should be on correctly specifying the deterministic trend component (up to its order), if this component is believed to be present.
A useful way to check the validity of the assumptions on the DGP is to consider placebo FAT estimates, obtained by pretending that the treatment was implemented before its actual implementation, and computing the corresponding placebo FAT. Placebo estimates that are statistically close to zero indicate that there is not enough evidence to reject the assumptions on the counterfactuals required by our method, [^9] That is, that the forecast of the counterfactual is unbiased on average, see Theorems 1 or 2. and give reassurance that FAT estimates the causal effect of the treatment (e.g., Arkhangelsky and Imbens (2023)).
Basis function regressions are easy to implement. Letting $i$ index an individual, the order of the basis function, $q_{i}$, and the length of the estimation window, $R_{i}$, are tuning parameters that must be chosen for each individual. If feasible, plotting the individual time series of pre-treatment data can guide this choice. If the panel is balanced (and the same basis function is used across individuals), one can simply plot the time series of pre-treatment outcomes averaged across individuals. If there are trends, one can decide if they are plausibly stochastic or deterministic. In the former case, any $q_{i}$ is valid, see Theorem 1. In the latter case, our result in Theorem 2 requires that $q_{i}$ be greater than or equal to the true order. A lower $q_{i}$ delivers smaller standard errors for FAT, while a higher $q_{i}$ could guard against nonstationary initial conditions (as we show in a simulation), which are otherwise ruled out by our class of DGPs. In practice, one could report results for a small range of values for $q_{i}$. The choice of estimation window $R_{i}$ depends on balancing a desire for small standard errors (larger $R_{i}$) with concerns about possible structural instability pre-treatment (smaller $R_{i}$). For example, assuming the same choice of tuning parameters across individuals, one can report FAT based on $q\in\{0,1,2\}$ if there are no obvious trends in averaged pre-treatment data and using a small $R$ (the smallest being $R=q+1$) if there are concerns about structural instability. Otherwise $R$ is all available pre-treatment data.
The basis function regression approach only uses pre-treatment outcomes as data and does not require specifying a model. We show that covariates, including lagged outcomes, can be incorporated in the estimation of FAT by specifying a model for the counterfactuals and using it to forecast the counterfactuals. We call this approach model-based FAT. This approach requires stronger assumptions, including: correct specification of the model; availability of a consistent estimator for the coefficients if these coefficients are homogeneous; symmetry of the error term if the coefficients are heterogeneous and the covariates are lagged outcomes. Under these additional assumptions, the model-based FAT is biased, but it is still consistent and asymptotically normal. A simulation illustrates how the model-based FAT is, however, sensitive to misspecification as well as estimation bias under correct specification in small samples. Estimation bias could, in principle, be corrected, but this requires strong assumptions on the DGP and, unlike the basis function approach, rules out unit roots. This suggests that the basis function approach has desirable robustness properties, particularly in small samples. Importantly, the basis function approach bypasses the incidental parameter problem in short panels.
The paper is related to several literatures. A growing body of work uses Bayesian or machine learning methods for forecasting counterfactuals from pre-treatment data of the treated (e.g., Brodersen et al. (2015); Chernozhukov et al. (2021); Menchetti et al. (2023); Cerqua et al. (2023)), while interrupted time series analysis (e.g., Bernal et al. (2017)) and “event studies” in finance (e.g., Brown and Warner (1985)) use a special case of the approach here to forecast counterfactuals. This work focuses on forecast accuracy [^10] See also the econometrics literature on forecasting with panel data, Baltagi (2013); Liu et al. (2020) and, typically, leverages the availability of both long time series and consensus about the plausible DGP for the counterfactuals. In contrast, we show that a long pre-treatment time series is not necessary for consistent estimation of the ATT. In fact, our approach is well-suited for short panels. We emphasize that a sufficient condition for consistent estimation of the ATT is forecast unbiasedness and show that basis function regressions are a robust method for obtaining unbiased forecasts under weak assumptions on the class of DGPs. Model-based approaches, instead, require stronger assumptions and can suffer from misspecification and/or estimation biases that translate into poor finite sample performance.
Forecast unbiasedness has been studied in the context of time series models (Fuller and Hasza (1980); Dufour (1984)), while with a short panel, the MLE-based method in Mavroeidis et al. (2015) could be used to produce average unbiased forecasts. Key assumptions in this work are a symmetry restriction on the model errors and stationarity. In this paper, we obtain unbiased forecasts for a more general class, including stationary and nonstationary processes with heterogeneous parameters. To the best of our knowledge, the issue of unbiased forecasts with a short panel and parameter heterogeneity has not been considered for non-stationary processes.
Our analysis allows for heterogeneous treatment effects. It is well known that in this case Ordinary Least Squares or Two Way Fixed Effects estimators in linear panel data models are generally inconsistent for the average treatment effect. [^11] See, Wooldridge (2005); Chernozhukov et al. (2013); de Chaisemartin and D’Haultfoeuille (2020). Proposed solutions assume the existence of a control group in every period. [^12] In Section C.2 we elaborate on the literature on program evaluation with control units. In addition, in order to avoid the incidental parameter problem, this literature assumes static panel models, i.e. the absence of lagged outcomes when the panel is short and the model allows for fixed effects, e.g., Angrist and Pischke (2009). Instead, our approach delivers a consistent estimator of the ATT in the absence of a control group and our class of DGPs allows for fixed effects, lagged outcomes, and time trends with fully heterogeneous parameters even when the panel is short.
Least-squares estimates of treatment effects are often interpreted as consistent estimates of certain weighted average effects to allow for the possibility of model misspecification, e.g., Theorem 1 in Chernozhukov et al. (2013). One generally has no control over the weights, which can become negative in panel data models with fixed effects. In addition, the weighted average interpretation does not address the incidental parameter problem, which occurs even for correctly specified models with homogeneous treatment effects. In contrast, our results show that, if treatment effects are estimated via averages over unbiased forecasts of counterfactuals, the correct unweighted treatment effect is consistently estimated, as long as the average is over a sufficiently large cross-section of observations. Our method thus signals a shift in focus: rather than trying to obtain consistent estimates of the model’s parameters, we focus on unbiased forecasts of counterfactuals. In this way, we avoid both the incidental parameter problem and the negative weighting issue mentioned above.
The paper is organized as follows: Section 2 presents the baseline method and Section 3 shows how to incorporate covariates. The finite sample performance of both methods is investigated in Section 4. We showcase the replication of an existing application without utilizing a control group in Section 5. Appendix A contains proofs and Appendix B discusses how to estimate the variance of our proposed estimator. An Online Appendix discusses extensions, including the existence of a control group, additional simulations and replications.
2 Baseline case: no control group, no covariates
In this section, we first show that our estimator is consistent and asymptotically normal under a high-level unbiasedness assumption for the forecasts of the counterfactuals. We then derive sufficient conditions on the class of DGPs and the forecast methods that satisfy the assumption.
@@ 2.1 Parameter of interest and estimator
Consider a treatment or a policy that is implemented at a time $\tau$. Here, the treatment affects all individuals in the population at the same time, [^13] In Section 2.4 we consider the more general case of staggered adoption, i.e., heterogeneous treatment timing across individuals. so that the treatment indicator of individual $i$ at time $t$ is given by
@@
We adopt the potential outcomes framework with each individual $i$ having two potential outcomes at each time $t$: $y_{it}(1)$ if the individual is exposed to the treatment and $y_{it}(0)$ if the individual is not exposed to the treatment. Due to the absence of a control group in our setting, we will henceforth simply refer to $y_{it}(0)$ as the “counterfactual”.
@@
$y_{it}$ =$=\left\{\begin{array}[]{ll}y_{it}(0)&\text{for}\;t\leq\tau,\\ y_{it}(1)&\text{for}\;t>\tau.\end{array}\right.$ (3)
@@
[…]
The challenge in identifying and estimating ${\rm ATT}_{h}$ is that the counterfactual $y_{i\tau+h}\left(0\right)$ is not observed for $h\geq 1$. The conventional approach in the presence of a control group is to impose sufficient assumptions that identify the parameter of interest from the observed post-treatment outcomes of the control group. In the absence of a control group, we exploit pre-treatment individual time series to obtain a forecast for $y_{i\tau+h}\left(0\right)$. We denote this forecast by $\widehat{y}_{i\tau+h}(0)$.
@@
For example, when $\{u_{i\tau+h}\}$ is a sequence of independent but not identically distributed random variables, Theorem 5.11 in White (2001) gives an asymptotic normality result.
Note that Assumption 2 allows for weak cross-sectional dependence. A key implication of the assumption is that it excludes shocks that affect all individuals after treatment and that are unforecastable. In principle, the assumption allows for some common shocks to be captured by the method used to forecast the counterfactuals. In the Online Appendix, we discuss how this assumption could be weakened in the presence of either an outcome variable subject to the same unforecastable shock but not subject to the treatment (Section C.1) or a control group subject to the same unforecastable shock (Section C.2).

### v4

@@ Abstract
We consider estimation and inference of the effects of a policy in the absence of an untreated or control group. We obtain unbiased estimators of individual (heterogeneous) treatment effects and a consistent and asymptotically normal estimator of the average treatment effect. Our estimator averages, across individuals, the difference between observed post‑treatment outcomes and unbiased forecasts of their counterfactuals, based on a (short) time series of pre-treatment data. The paper emphasizes the importance of focusing on forecast unbiasedness rather than accuracy when the end goal is estimation of average treatment effects. We show that simple basis function regressions ensure forecast unbiasedness for a broad class of data generating processes for the counterfactuals. In contrast, forecasting based on a specific parametric model requires stronger assumptions and is prone to misspecification and estimation bias. We show that our method can replicate the findings of some previous empirical studies but it does so without using an untreated or control group.
@@ 1 Introduction
Evaluating the effect of a policy or a treatment usually requires comparing treated and untreated units. Standard approaches such as difference-in-differences (DiD) or two-way fixed effects (TWFE) regressions rely on outcomes of untreated units to identify the counterfactual outcomes of treated units. These methods rely on assumptions such as unconfoundedness or parallel trends (e.g., Goodman-Bacon 2021; Callaway and Sant’Anna 2021; Sun and Abraham 2020). They cannot be applied when all units receive the treatment.
This situation is common. Examples include national tax reforms, health programs, and environmental regulations that affect all individuals or regions at the same time. Even when adoption is staggered, the last periods after full treatment do not contain untreated units, so conventional DiD or event-study designs cannot recover treatment effects for those periods. In such cases, researchers often turn to structural models or to simple before–after comparisons. These settings call for methods that make the required assumptions explicit and allow researchers to evaluate their plausibility.
This paper studies how to estimate treatment effects when no valid control group exists. We propose a transparent and practical approach based on forecasting each treated unit’s counterfactual outcome from its own pre-treatment data using basis function regressions. The key idea is that in short panels, the property required for identification is forecast unbiasedness rather than forecast accuracy. Forecasting methods in time-series and machine learning aim to minimize prediction error and often allow some bias to reduce variance. In policy evaluation, this objective is not appropriate. An unbiased forecast of the counterfactual outcome guarantees an unbiased estimate of the treatment effect, while a forecast that is accurate but biased does not. Of course, our approach relies on its own identifying assumptions, which we make explicit, but they differ from those used in designs that compare treated and untreated units.
Our approach, the Forecasted Average Treatment effect (FAT) estimator, computes the average treatment effect on the treated as the average difference between the observed post-treatment outcome and an unbiased forecast of the outcome in the absence of treatment. What matters is that the forecasting rule produces unbiased predictions on average across individuals.
The paper makes three contributions. First, we develop a simple and general framework for estimating treatment effects when no untreated units are available. Under appropriate assumptions, the FAT estimator is consistent and asymptotically normal even when the time dimension is short, the panel is unbalanced, and treatment effects are heterogeneous.
Second, we characterize a broad class of data generating processes (DGPs) under which the required forecast unbiasedness condition holds. When individual counterfactual outcomes can be generally expressed as the sum of (at most) three unobserved components - a stationary process, a unit root process, and a deterministic trend - unbiased forecasts can be obtained by regressing pre-treatment outcomes on basis functions of time, such as low-order polynomials. This result allows for DGPs with fixed effects, heterogeneous autoregressive parameters, and non-stationary trends, without requiring a detailed specification of the stochastic component. The framework also accommodates heterogeneous dynamic processes, such as unit-specific autoregressive parameters, which are difficult to handle with standard short-panel methods. We call our proposed approach “Unobserved-Components FAT”.
Third, we compare our general “Unobserved-Components FAT” approach with what we refer to as “Parametric-Model FAT”, where the researcher specifies and estimates a particular parametric model such as an AR(1) to forecast counterfactuals. This approach relies on stronger assumptions and can perform poorly in short panels because of misspecification or estimation bias. In contrast, our regression-based approach that focuses directly on forecast unbiasedness is robust to specification errors and is not affected by the incidental parameter problem.
The FAT estimator can also be used when untreated units exist but are imperfect comparators. In such cases, FAT can provide a robustness check for studies that rely on, e.g., parallel-trends assumptions. The framework complements recent work on DiD and event-study designs (e.g., de Chaisemartin and D’Haultfoeuille 2020; Callaway and Sant’Anna 2021; Sun and Abraham 2020) by providing an alternative benchmark based on explicit assumptions about the individual-specific counterfactual process rather than on cross-group comparability.
Crucially, the approach based on FAT is advantageous when outcomes exhibit dynamics. Recent work (Marx et al., 2024; Klosin, 2024; Botosaru and Liu, 2025; Cornwall and Wentland, 2025) has shown that traditional TWFE and dynamic panel estimators often yield biased or negatively weighted estimates of the treatment effect in the presence of serial correlation. The solutions proposed in this literature typically rely on the homogeneity of the persistence parameters across units. By construction, FAT avoids conflating structural treatment effects with outcome persistence and is robust to feedback mechanisms (Bonhomme, 2025), while allowing for a much richer class of models, which include heterogeneous forms of serial correlation, where the degree of outcome persistence can vary arbitrarily across individuals.
This paper relates to several strands of literature. It is thematically related to work on causal inference in settings where contemporaneous untreated comparison groups are not the primary source of identification, and where time-series structure plays a central role in constructing counterfactual outcomes from pre-treatment data. These include Bayesian approaches to causal forecasting (e.g., Brodersen et al. 2015); (comparative) interrupted time-series methods, in which untreated potential outcomes are modeled using finite-dimensional parametric mean functions in time—typically linear or piecewise-linear trends with level and/or slope changes at the intervention date (e.g., Schochet 2022; Bernal et al. 2017; Brown and Warner 1985); and methods based on data-driven aggregation to purge unobserved aggregate confounders in designs with aggregate shocks (e.g., Arkhangelsky and Korovkin 2024).
The FAT framework differs from these strands along three dimensions. First, relative to Bayesian causal forecasting, FAT is designed to deliver forecast-unbiased counterfactuals under high-level restrictions on the untreated outcome process and does not require the researcher to commit to a fully specified likelihood or prior. Second, relative to interrupted time-series methods, FAT strictly nests the canonical segmented-regression mean specifications while allowing for richer latent evolution, including stochastic trend components and heterogeneous unit-specific dynamics. Third, relative to the aggregate-shock/aggregation literature, FAT does not take identification from exogenous aggregate shocks or instruments, nor does it rely on cross-sectional weighting to eliminate unobserved aggregate confounding; instead, it is a short-panel forecasting approach in which identification is driven by the internal time-series information in treated units’ pre-intervention histories.
In addition, the paper contributes to the panel-data literature on heterogeneous dynamic processes and random coefficients (e.g., Chamberlain 1992; Arellano and Bonhomme 2012; Graham and Powell 2012). For example, the projection estimator in Arellano and Bonhomme (2012) can be interpreted as a special case of FAT under specific restrictions on the latent outcome process. In particular, for a unit-root process without drift, the Arellano–Bonhomme forecast corresponds to an equal-weighted average of pre-treatment outcomes, whereas the FAT polynomial forecast assigns greater weight to more recent observations. When deterministic trends are present, the projection approach in Arellano and Bonhomme (2012) requires correct specification of the trend component, while FAT only requires the choice of a sufficiently rich family of basis functions, such as low-order polynomials. Therefore, FAT delivers unbiased counterfactual forecasts under weaker restrictions on the trend component, while retaining the projection-based intuition underlying existing random-coefficient panel estimators.
Our analysis also relates to work on forecast unbiasedness in time-series models (e.g., Fuller and Hasza 1980; Dufour 1984) and to the dynamic-panel literature on bias correction and consistent estimation with short time dimensions (e.g., Angrist and Pischke 2009; Blundell and Bond 1998). Our approach differs by showing that consistent treatment effect estimation in short panels can be achieved through unbiased forecasts rather than through the specification of a full stochastic model.
The rest of the paper is organized as follows. Section 2 presents the framework and the FAT estimator. Section 3 extends the method to include covariates and discusses forecasting based on a parametric model. Section 4 reports simulation results. Section 5 provides an empirical illustration. Section 6 concludes.
2 Baseline case: Unobserved-Components FAT
In this section we develop our baseline “Unobserved-Components FAT” estimator for the average treatment effect on the treated (ATT) under universal treatment. We use individual pre‑treatment time series to forecast each unit’s post‑treatment counterfactual outcome and then average, across units, the difference between the observed outcome and its forecast. We first establish consistency and asymptotic normality of this estimator under a high‑level forecast‑unbiasedness condition, and then show that this condition holds for a broad class of unobserved‑components DGPs when counterfactuals are forecast using basis‑function regressions.
@@ 2.1 Parameter of interest and estimator
Consider a treatment or a policy that is implemented at a time $\tau$. [^3] In Section 2.4, we discuss individual-specific timing. Here, the treatment affects all individuals in the population at the same time, so that the treatment indicator of individual $i$ at time $t$ is given by
@@
We adopt the potential outcomes framework with each individual $i$ having two potential outcomes at each time $t$: $y_{it}(1)$ if the individual is exposed to the treatment and $y_{it}(0)$ if the individual is not exposed to the treatment. Due to the absence of an untreated group in our setting, we will henceforth simply refer to $y_{it}(0)$ as the “counterfactual".
@@
$y_{it}$ $=\left\{\begin{array}[]{ll}y_{it}(0)&\text{for}\;t\leq\tau,\\ y_{it}(1)&\text{for}\;t>\tau.\end{array}\right$. (3)
@@
where we used that $y_{i\tau+h}(1)=y_{i\tau+h}$ for $h\geq 1$. [^4] Note that with identically and independently distributed data across $i$, the right hand side of (4) reduces to the conventional $\mathbb{E}\left[y_{i\tau+h}-y_{i\tau+h}\left(0\right)\right]$.
The challenge in identifying and estimating ${\rm ATT}_{h}$ is that the counterfactual $y_{i\tau+h}\left(0\right)$ is not observed for $h\geq 1$. The conventional approach in the presence of an untreated group is to impose sufficient assumptions that identify the parameter of interest from the observed post-treatment outcomes of the untreated group. In the absence of an untreated group, we exploit pre-treatment individual time series to obtain a forecast for $y_{i\tau+h}\left(0\right)$. We denote this forecast by $\widehat{y}_{i\tau+h}(0)$.
@@
For example, when $\{u_{i\tau+h}\}$ is a sequence of cross-sectionally independent (but not identically distributed) random variables, Theorem 5.11 in White (2001) gives an asymptotic normality result.
What Assumptions 1 and 2 fundamentally rule out is shocks that affect all individuals after the treatment and that are unforecastable. [^7] In principle, the assumptions allow for some common shocks to be captured by the method used to forecast the counterfactuals. For example, if the common shocks are deterministic and can be modeled as polynomials, our baseline method will be able to account for it. This is an example of the assumptions that any method for treatment effect evaluation must inevitably make in the absence of a valid control group. The assumptions are more likely to be valid the higher the frequency at which the data is measured and the closer to the treatment date one evaluates the effect (i.e., they more plausibly hold for monthly data and for effects evaluated one month after the treatment than for yearly data and effects evaluated several years after the treatment). Below we provide some discussion of how the assumption of no common shocks after the treatment could be weakened. As one would expect, the assumption can be relaxed when there is a group of untreated individuals that are subject to the same unforecastable shock, as we discuss in the Online Appendix (Section C.1). Perhaps less obviously, the assumption could also be relaxed in the absence of an untreated group, as long as we observe another variable for the treated units that is subject to the same unforecastable shock but not subject to the treatment (see the discussion in Section 6).